OpenStudy (azionne):

15(2 - x) = 13(3 - x) will medal

OpenStudy (skullpatrol):

What does the question ask for? :-)

OpenStudy (skullpatrol):

Are you still there?

OpenStudy (azionne):

hold on

OpenStudy (skullpatrol):

np

OpenStudy (azionne):

a reduced mixed number

OpenStudy (skullpatrol):

For what?

OpenStudy (azionne):

for the equation above

OpenStudy (skullpatrol):

Ok, so that means they want you to solve for x, right?

OpenStudy (azionne):

yes

OpenStudy (skullpatrol):

Do you recall the distributive property?

OpenStudy (azionne):

No

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (azionne):

lol do u mind just giving me this one

OpenStudy (skullpatrol):

See how the "a" distributes over the "b" and the "c"? This is important to really learn and understand :-)

OpenStudy (azionne):

yes i know not right now though if i get this i will graduate highschool

OpenStudy (skullpatrol):

distribute the "15" and the "13" please

OpenStudy (azionne):

30-15x=39-13x

OpenStudy (skullpatrol):

Excellent! Good work. Now, subtract 30 from BOTH sides.

OpenStudy (azionne):

-15x=9-13x

OpenStudy (skullpatrol):

Next, add 13x to BOTH sides.

OpenStudy (azionne):

-2x=9

OpenStudy (skullpatrol):

what do you think you should do next? :-)

OpenStudy (skullpatrol):

recall you want "x" alone on one side

OpenStudy (azionne):

x=-0.2 repeated

OpenStudy (skullpatrol):

why?

OpenStudy (azionne):

idk this is what i got

OpenStudy (skullpatrol):

In your equation: $$\Huge-2x=9$$ divide BOTH sides by -2

OpenStudy (azionne):

i did i already gave u my answer

OpenStudy (skullpatrol):

how did you calculate -0.2 repeated?

OpenStudy (azionne):

idk

OpenStudy (skullpatrol):

did you guess? :-D

OpenStudy (skullpatrol):